### v1

@@ 6 Discussion and conclusions
The measurements of boron and nitrogen surface abundances in main sequence B type stars of this young open star cluster provide a unique opportunity to shed light on the origin of contamination of a significant fraction of these stars with the products of hydrogen fusion. The best quantitative interpretation of these measurements in terms of rotationally induced mixing in the stellar interior requires us to disregard the five most slowly rotating of our targets, which we tentatively consider as binary products where mixing mechanisms other than rotation dominate. We can then conclude that rotational mixing appears likely as explanation of the observed abundances of the remaining stars, since the trends of the surface abundances with stellar rotation and luminosity are in good agreement with the observations. Obtaining the best quantitative fit to the data does require a reduced mixing efficiency, , to best fit the data, which is significantly lower than either the value of 0.022 adopted by Chaboyer & Zahn (1992); Heger et al. (2000); Brott et al. (2011a), or the standard value of 0.033 adopted by the models of Frischknecht et al. (2010).
@@
The binary interpretation of the slow rotators is only based on circumstantial evidence (cf. Sect.5). A rigorous analysis including the surface abundances of the slow rotators must await corresponding models predictions (Jin et al., in prep.). Nevertheless, we were able to construct a working hypothesis that the majority of rapidly rotating upper main sequence stars undergo mild rotationally induced mixing while the slow rotators are mostly merger products. While this is free of immediate contradictions, it is based on only 18 total stars, and so our conclusions may be sensitive to statistical fluctuations. It also leaves out the fastest rotators in young open clusters, i.e., the Be stars, which have been suggested to originate from mass transfer in binaries (Pols et al. 1991 A&A, 241, 419; Hastings et al. 2021A&A…653A.144), and the roles of pulsations (Varghese et al., 2023ApJ…942…53) and magnetic fields (Meynet et al. 2011A&A…525L..11), which could amplify or reduce internal mixing. Further observations and stellar interior calculations are thus needed to come to more definite conclusions.
@@
Acknowledgements Based on observations with the NASA/ESA Hubble Space Telescope obtained [from the Data Archive] at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Incorporated, under NASA contract NAS5-26555. Support for program number 14673 was provided through a grant from the STScI under NASA contract NAS5-26555. Much of the data presented in this paper were obtained from the Mikulski Archive for Space Telescopes (MAST) at the Space Telescope Science Institute. The specific observations analyzed can be accessed via https://doi.org/10.17909/7vgt-zq5 (catalog https://doi.org/10.17909/7vgt-zq5). STScI is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS5–26555. Support to MAST for these data is provided by the NASA Office of Space Science via grant NAG5–7584 and by other grants and contracts. SD acknowledges CNPq/MCTI for grant 306859/2022-0. DJL is supported by the Spanish Government Ministerio de Ciencia e Innovación through grants PID2021-122397NB-C21 and SEV 2015-0548. This paper also made use of data obtained from the ESO Science Archive Facility and observed using ESO Telescopes at the La Silla Paranal Observatory under ESO observing program IDs 73.D-0234, 171.D-0237, 081.A-9006, 081.A-9008, and 188.B-3002. We also thank Jesús Maíz Apellániz for useful discussions on the distance to NGC 3293.

### v2

@@ 6 Discussion and conclusions
The measurements of boron and nitrogen surface abundances in main sequence B type stars of this young open star cluster provide a unique opportunity to shed light on the origin of contamination of a significant fraction of these stars with the products of hydrogen fusion. The best quantitative interpretation of these measurements in terms of rotationally induced mixing in the stellar interior requires us to disregard the five most slowly rotating of our targets, which we tentatively consider as binary products where mixing mechanisms other than rotation dominate. We can then conclude that rotational mixing appears likely as explanation of the observed abundances of the remaining stars, since the trends of the surface abundances with stellar rotation and luminosity are in good agreement with the observations. We find the boron data is better reproduced with a rotational mixing efficiency parameter which is reduced by a factor of two (), with respect to the widely adopted value of .
@@
The binary interpretation of the slow rotators is only based on circumstantial evidence (cf. Sect.5). A rigorous analysis including the surface abundances of the slow rotators must await corresponding models predictions (Jin et al., in prep.). Nevertheless, we were able to construct a working hypothesis that the majority of rapidly rotating upper main sequence stars undergo mild rotationally induced mixing while the slow rotators are mostly merger products. While this is free of immediate contradictions, it is based on only 18 total stars, and so our conclusions may be sensitive to statistical fluctuations. It also leaves out the fastest rotators in young open clusters, i.e., the Be stars, which have been suggested to originate from mass transfer in binaries (Pols et al., 1991; Hastings et al., 2021), and the roles of gravity waves (Varghese et al., 2023) and magnetic fields (Meynet et al., 2011), which could amplify or reduce internal mixing. Further observations and stellar interior calculations are thus needed to come to more definite conclusions.
@@
Acknowledgements Based on observations with the NASA/ESA Hubble Space Telescope obtained [from the Data Archive] at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Incorporated, under NASA contract NAS5-26555. Support for program number 14673 was provided through a grant from the STScI under NASA contract NAS5-26555. Much of the data presented in this paper were obtained from the Mikulski Archive for Space Telescopes (MAST) at the Space Telescope Science Institute. The specific observations analyzed can be accessed via https://doi.org/10.17909/7vgt-zq51 (catalog https://doi.org/10.17909/7vgt-zq51). STScI is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS5–26555. Support to MAST for these data is provided by the NASA Office of Space Science via grant NAG5–7584 and by other grants and contracts. SD acknowledges CNPq/MCTI for grant 306859/2022-0. DJL is supported by the Spanish Government Ministerio de Ciencia e Innovación through grants PID2021-122397NB-C21 and SEV 2015-0548. HJ received financial support for this research from the International Max Planck Research School (IMPRS) for Astronomy and Astrophysics at the Universities of Bonn and Cologne. The authors gratefully acknowledge the granted access to the Bonna cluster hosted by the University of Bonn. This paper also made use of data obtained from the ESO Science Archive Facility and observed using ESO Telescopes at the La Silla Paranal Observatory under ESO observing program IDs 73.D-0234, 171.D-0237, 081.A-9006, 081.A-9008, and 188.B-3002. We want to thank the anonymous referee for providing a careful review and insightful comments that resulted in significant improvements to the final paper. We also thank Jesús Maíz Apellániz for useful discussions on the distance to NGC 3293.